Circular No. 7217

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

GRB 990704 J. Heise and J. in ’t Zand, Space Research Organization of the Nether-

lands (SRON), Utrecht; G. Celidonio, G. Scotti, and G. Gandolfi, BeppoSAX Science Operation Center, Telespazio, Rome; M. Feroci, P. Soffitta, and L. Piro, Istituto di Astrofisica Spaziale (IAS), CNR, Rome; and E. Palazzi and L. Amati, Istituto Tecnologie e Studio Radiazioni Extraterrestri (ITESRE), Bologna, report: “A γ-ray burst with a bright x-ray counterpart triggered the BeppoSAX Gamma-Ray Burst Monitor (GRBM) on July 4.7294 UT and was simultaneously detected in the Wide Field Camera (WFC) unit 1. The burst lasted for ∼ 40 s in the 2–26-keV WFC data and had three peaks with increasing intensity. The WFC peak intensity corresponds to a flux of 6.2 Crab. The 40–700-keV GRBM flux is weak and has a similar duration. A peak of 320 counts/s is coincident with the maximum intensity in the WFC. The WFC-determined position and error circle of the x-ray counterpart to GRB 990704 is as noted on IAUC 7214 (99-percent confidence level), which includes a potential 4′ error in the satellite attitude.”

M. Feroci and G. Gandolfi, IAS, Rome; L. A. Antonelli, Osservatorio Astronomico di Roma; M. R. Daniele, C. De Libero, L. Di Ciolo, and V. Torroni, BeppoSAX, Rome; F. Frontera, ITESRE, CNR, Bologna; and J. Heise, SRON, report on behalf of the BeppoSAX Team: “The BeppoSAX Narrow Field Instruments observed the region of GRB 990704 from July 5.064 to 5.989 UT. One previously unknown x-ray source (1SAX J1219.5–0350) was found inside the 7′ BeppoSAX WFC error circle at α = 12^{h}19^{m}27^{s}.3, δ = −3^{o}50′22′′ (equinox 2000.0, error radius 1′, 90-percent confidence level). The average source flux at 2–10 keV was (8.0 ± 0.6) × 10^{−3} counts/s, corresponding to ∼ 4.4 × 10^{−13} erg cm^{−2} s^{−1}, assuming a power-law spectrum. The source shows a fading behavior, strongly suggesting that it is the x-ray afterglow of GRB 990704.”

Regarding their report on IAUC 7214, A. Maury et al. write: “We have observed the field of GRB 990704 again during the night of July 5–6. The optical object reported on IAUC 7214 is still visible at the same position and seems to be not variable (see also GCN 367, 370, 371); it is therefore very likely not the optical counterpart of GRB 990704.”

1999 July 6  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green